Show that for an ideal gas Cp=yR/(y-1), Cv=R/(Y-1) and Cp-Cv=R
Answer #1

Please find the attachment.prove [cp-CV=R [cv = 8/12-1)] [cp = YR/(1-1) де кто о #ket, HEU + PV foạide91 995 je pv=nRT H=ut MRT Whene, Differenciate botput cp = Y Cv in equation ③ anny y cy - Cv=R CV (Y-1) =R [cv = 2 ] -